Verify that y = e2xsinx is a solution to the differential equation d2y/dx2 - 4dy/dx+5y=0
1
Expert's answer
2020-06-30T17:53:32-0400

If "y=e^{2x}\sin x" then "y'=e^{2x}(2\sin x+\cos x)" and "y''=2e^{2x}(2\sin x+\cos x)+e^{2x}(2\cos x-sin x)" hence "y''-4y'+5y=0"